OpenStudy (anonymous):

Solve for x. 7(x - 3) + 3(4 - x) = -8

OpenStudy (uri):

7x -21+12-3x=-8 4x-9+8=0 4x-1=0

OpenStudy (imtiaz7):

then 4x-1=0 4x=1 x=1/4

OpenStudy (anonymous):

X=\[\frac{ 1 }{4 }\]